OpenStudy (anonymous):

I Need help on solving this arithmetic sequence...... a1=10, d=5; find a13

OpenStudy (diyadiya):

\[a_n= a_1 +(n-1)d\]\[\Large a_{13}=a_1 + (13-1)d=a_1 + 12d\]

OpenStudy (anonymous):

What about the d=5 part . or is that the answer ??

OpenStudy (diyadiya):

You have to substitute the value in a1+ 12d ( a1=10 ,d=5)

OpenStudy (anonymous):

thank you

OpenStudy (diyadiya):

Anytime :) so what is your answer?

OpenStudy (anonymous):

idk im horrible at math ...

OpenStudy (diyadiya):

a13=a1+12d =10 + (12 x 5) = ?

OpenStudy (anonymous):

70?

OpenStudy (diyadiya):

Yes :) Remember a_n= a1 + (n-1)d

OpenStudy (anonymous):

thank you..

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

So a13 = 70?

OpenStudy (diyadiya):

Yes

OpenStudy (anonymous):

ok what about this one? : a1=6, d=9 find a8

OpenStudy (anonymous):

a1=6, d=9 find a8

OpenStudy (diyadiya):

so what is the equation ,i just gave you that in the beginning

OpenStudy (anonymous):

idk ... =\ . pleasee helpppp!

OpenStudy (diyadiya):

\[a_n=a_1 + (n-1)d\] here n is 8

OpenStudy (anonymous):

an=a1+(8−1)d

OpenStudy (diyadiya):

yes!! so ...

OpenStudy (diyadiya):

a8=a1+(8−1)d

OpenStudy (anonymous):

Need some more help. ..! =\ don't remember what's next.

OpenStudy (diyadiya):

a8=a1+(8−1)d =a1+7d

OpenStudy (anonymous):

a8=8 ???

OpenStudy (diyadiya):

how?

OpenStudy (anonymous):

1+7=8

OpenStudy (diyadiya):

Noo what is a1 & d given in question ? a8=a1 + 7(d)

OpenStudy (anonymous):

i don't know... =\ hmmph , i feel stupid .

OpenStudy (diyadiya):